OpenStudy (anonymous):

Pleasee Helpp? D: Given the function f(x) = log2(x + 6), find the value of f^−1(3).

random231 (random231):

take y=log2(x+6)

hartnn (hartnn):

have you tried this ?

OpenStudy (anonymous):

I got f^−1(3) = 9 when I tried, but idk if it's correct

hartnn (hartnn):

and how ? no its not correct.

hartnn (hartnn):

\(3 = \log_2(x+6) \implies x+6 = 3^2\) can you get x from there ?

OpenStudy (anonymous):

I separated the equation and got f(x) =2log(X)+2log(6), but then didn't know what to do so I guessed by looking at the numbers. And yes I can get x now. I see what you did. Thank you!

OpenStudy (anonymous):

I got 3 @ hartnn

hartnn (hartnn):

3 is correct :)

hartnn (hartnn):

oh wait!

hartnn (hartnn):

\(3 = \log_2(x+6) \implies x+6 = 2^3\) sorry for that mistake

hartnn (hartnn):

x+6 =8 x=2

hartnn (hartnn):

@Jessicaa

OpenStudy (anonymous):

ohh okay. Thank you for the help!

hartnn (hartnn):

welcome ^_^

hartnn (hartnn):

since you seem to be new here, \(\Huge \mathcal{\text{Welcome To OpenStudy}\ddot\smile} \)

OpenStudy (anonymous):

lol thank you. I am. I'm still learning

hartnn (hartnn):

have fun!